Working Together, Beth and Anna can clean their apartment in three hours. If Ana had done it alone, it would have taken her seven hours. How long would it have taken Beth if she had worked alone.
Answer by stanbon(75887)

Working Together, Beth and Anna can clean their apartment in three hours. If Ana had done it alone, it would have taken her seven hours. How long would it have taken Beth if she had worked alone.
------
Equate rates::
b + a = 1/3
a = 1/7
b = (1/3)-(1/7)
b = (7-3)/21 = 4/21 (Beth's rate
----
Ans:: Beth's time = 21/4 = 5 hr 15 min (to do the job alone)
--------------
Cheers,
Stan H.
--------------